# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v2

@@ 2.2 The $e_{x}$, the Dirac-like inf-irreducible functions of $\mathcal{G}$
We now turn to our candidates for inf-irreducible elements of a complete subspace $\mathcal{G}$ of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$. Define the Archimedean class of a function $f\in\mathcal{G}$ as
$[f]:=\{g\in\mathcal{G}\,|\,\exists\alpha\,\in\mathbb{R},\,f-\alpha\leq g\leq f+\alpha\}.$ (17)
Let us put an order on Archimedean classes, saying that $[f]\leq[g]$ if there exists $\alpha\in\mathbb{R}$ such that $f\leq g+\alpha$. We say that $[f]$ is maximal if $[f]\leq[g]\implies[f]\geq[g]$.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. ¯$\overline{\mathbb{R}}^{\mathscr{Y}}$) for some sets $\mathscr{X}$ and $\mathscr{Y}$. Let $J$ be a (max,+)-isomorphism from $\mathscr{F}$ onto $\mathcal{G}$. If $f\in\mathscr{F}$ is such that $[f]$ is maximal, then $[Jf]$ is also maximal.
@@
Take $g_{0}\in\mathcal{G}$ and $\alpha\in\mathbb{R}$ such that $Jf\leq g_{0}+\alpha$. Fix $f_{0}\in\mathscr{F}$ such that $g_{0}=Jf_{0}$, then $f\leq f_{0}+\alpha$. Since $[f]$ is maximal, there exists $\alpha^{\prime}\in\mathbb{R}$, $f\geq f_{0}+\alpha^{\prime}$, composing with $J$, we obtain that $Jf\geq g_{0}+\alpha^{\prime}$, so $[Jf]=[g_{0}]$ and $[Jf]$ is maximal. ∎
It is thus tempting to focus on maximal Archimedean classes, and we will do so in 23. However some sets have no maximal classes as seen in 18.
We will often need in the sequel to consider at least some finite values, hence we restrict ourselves to proper sets $\mathcal{G}$ of functions of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$. We are also not interested in situations where points cannot be distinguished by $\mathcal{G}$, since we could then quotient $\mathscr{X}$ by a $\mathcal{G}$-induced equivalence relation. Hence we introduce
@@
Let $\mathcal{G}$ be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ for some set $\mathscr{X}$, with $\mathcal{G}$ satisfying 1. Define, for any $x\in\mathscr{X}$, the function $e_{x}:\mathscr{X}\rightarrow\overline{\mathbb{R}}$ by
@@
with $\infty$- ∞ absorbing. Moreover, for any $f\in\mathcal{G}$, we have the representation (with +$\infty$+ ∞ absorbing)
@@
and $e_{x}$ satisfies a triangular inequality: for all $x,x^{\prime},x^{\prime\prime}\in\mathscr{X}$, $e_{x}(x^{\prime\prime})\leq e_{x^{\prime}}(x^{\prime\prime})+e_{x}(x^{\prime})$. If $f\in\mathcal{G}$ is such that $[f]$ is maximal, then for all $x_{0}\in\operatorname*{Dom}(f)$ we have $[e_{x_{0}}]=[f]$, i.e. we can fix $\lambda_{0}\in\mathbb{R}$, such that $e_{x_{0}}+\lambda_{0}\leq f$. If $\mathcal{G}$ contains the constant functions, then $e_{x}$ has nonnegative values.
@@
Let $x\in\mathscr{X}$. Since $\mathcal{G}$ is proper, we can fix $u_{x}\in\mathcal{G}$ such that $u_{x}(x)\in\mathbb{R}$. So $e_{x}\geq u_{x}-u_{x}(x)$, whence $e_{x}(x)\geq 0$, while by definition $e_{x}(x)\leq 0$. We have that $e_{x}\in\mathcal{G}$ as $\mathcal{G}$ is sup-stable. If $e_{x}=\inf(f,g)$ for some $f,g\in\mathcal{G}$ then w.l.o.g. we can assume that $f(x)=e_{x}(x)=0$ so, by maximality of $e_{x}$, $f\leq e_{x}=\inf(f,g)$ and $f=e_{x}$. Take $f\in\mathcal{G}$ with $[f]$ maximal. Since $f\leq e_{x}+f(x)$ for all $x\in\operatorname*{Dom}(f)$ i.e. $[e_{x}]\geq[f]$, we have that $[e_{x}]\leq[f]$ implies that $[e_{x}]=[f]$. Fix $x\neq x^{\prime}$ and $\lambda\in\overline{\mathbb{R}}$ and $g_{1},g_{2}\in\mathcal{G}$ as in 1. Assume that $e_{x}=e_{x^{\prime}}+\lambda$. We can already exclude $\lambda\in\{\pm\infty\}$ since $e_{x}$ has a finite value in $x$. Moreover by definition of $e_{x}$ we have
@@
Evaluating in $x^{\prime}$ and $x$ the above expressions gives $\lambda\geq g_{1}(x^{\prime})-g_{1}(x)\geq\lambda$. The same holds for $g_{2}$, which contradicts 1. So $e_{x}\neq e_{x^{\prime}}+\lambda$. Eq.(19) is just a rewriting of (18).
Concerning (20), on the one hand, we can take the infimum over $x\in\mathscr{X}$ of $f\leq e_{x}+f(x)$, the latter giving the triangular inequality when taking $f=e_{x^{\prime}}$. On the other hand, for any $y\in\mathscr{X}$, we have that $\inf_{x\in\operatorname*{Dom}(f)}e_{x}(y)+f(x)\leq e_{y}(y)+f(y)=f(y)$. ∎
@@
Consider the set $\mathcal{G}$ of increasing functions from $\mathbb{R}$ to $\overline{\mathbb{R}}$. It is sup and inf-stable. A quick computation gives $e_{x}=\delta^{\top}_{\{y\in\mathbb{R}|y\leq x\}}$ where $\delta^{\top}_{A}$ denotes the indicator function of a subset $A\subset\mathbb{R}$. This $e_{x}$ looks like a Dirac function, but $[e_{x}]$ is strictly smaller than $[e_{x^{\prime}}]$ for any $x^{\prime}<x$, so no Archimedean class is maximal. Moreover the constant functions are inf-irreducible, but do not belong to $\{e_{x}+\lambda\}_{\lambda\in\mathbb{R},x\in\mathbb{R}}$.
@@
Consider the inf-closure $\overline{\mathcal{G}}^{\inf}$ of a complete subspace $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ satisfying 1. Then any $f\in\mathcal{G}$ such that $f$ is inf-irreducible in $\overline{\mathcal{G}}^{\inf}$ is inf-irreducible in $\mathcal{G}$ and
@@
Set $\tilde{e}_{x}^{<}(\cdot)=\sup\{u\in\overline{\mathcal{G}}^{\inf}\,|\,u(x)<0\}$ and $\tilde{e}_{x}(\cdot)=\sup\{u\in\overline{\mathcal{G}}^{\inf}\,|\,u(x)\leq 0\}$. By definition, $\tilde{e}_{x}^{<}\leq\tilde{e}_{x}$, and, for all $\epsilon>0$, $\tilde{e}_{x}-\epsilon\leq\tilde{e}_{x}^{<}$. Taking the limit $\epsilon\to 0^{+}$, $\tilde{e}_{x}^{<}=\tilde{e}_{x}$. Since $\mathcal{G}\subset\overline{\mathcal{G}}^{\inf}$, $e_{x}\leq\tilde{e}_{x}$. Let $u=\inf_{\alpha\in\mathcal{A}}h_{\alpha}\in\overline{\mathcal{G}}^{\inf}$ such that $u(x)<0$ and fix $\alpha\in\mathcal{A}$ such that $h_{\alpha}(x)\leq 0$. As $h_{\alpha}\in\mathcal{G}$, we have $u\leq h_{\alpha}\leq e_{x}$. Taking the supremum over such $u$ entails that $\tilde{e}_{x}^{<}\leq e_{x}$, which concludes the proof. ∎
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. ¯$\overline{\mathbb{R}}^{\mathscr{Y}}$) for some sets $\mathscr{X}$ and $\mathscr{Y}$, with both $\mathscr{F}$ and $\mathcal{G}$ satisfying 1. Let $J$ be a (max,+)-isomorphism from $\mathscr{F}$ onto $\mathcal{G}$ and define $e_{x}\in\mathcal{G}$ (resp. $e^{\prime}_{y}\in\mathscr{F}$) as in (18). Then the following statements are equivalent:
@@
    $J(\{e^{\prime}_{y}+\lambda^{\prime}\}_{y\in\mathscr{Y},\,\lambda^{\prime}\in\overline{\mathbb{R}}})=\{e_{x}+\lambda\}_{x\in\mathscr{X},\,\lambda\in\overline{\mathbb{R}}}$;
@@
Fix $y\in\mathscr{Y}$ and recall that $e^{\prime}_{y}(\cdot):=\sup\{v\in\mathscr{F}\,|\,v(y)\leq 0\}$. By assumption i), $Je^{\prime}_{y}=e_{\psi(y)}+\bar{g}(y)$ with $\bar{g}(y)=Je^{\prime}_{y}(\psi(y))\in\mathbb{R}$. Conversely there exists a $\varphi:\mathscr{X}\to\mathscr{Y}$ and $\lambda_{x}\in\mathbb{R}$ such that $e_{x}=J(e_{\varphi(x)}+\lambda_{x})=e_{\psi(\varphi(x))}+\lambda_{x}+\bar{g}(\varphi(x))$. However from 17, we know that 1 then ensures that $x=\psi(\varphi(x))$, so $\psi$ is surjective. Similarly, consider $Je^{\prime}_{y}=Je^{\prime}_{y^{\prime}}$, to get that $\psi$ is injective.
@@
so i) holds. Assume that $Jf=g+f\circ\phi=\tilde{g}+f\circ\tilde{\phi}$ for some alternative $\tilde{g}$ and $\tilde{\phi}$ satisfying (23). Then the above computation gives that $g(\phi^{-1}(y))+e_{\phi^{-1}(y)}=\tilde{g}(\tilde{\phi}^{-1}(y))+e_{\tilde{\phi}^{-1}(y)}$ for all $y\in\mathscr{Y}$. 17 then gives that necessarily $\phi=\tilde{\phi}$, and, evaluating in $\phi^{-1}(y)$, $g=\tilde{g}$. So the representation (23) is unique.
@@
$Ju(x)-u(\phi(x))+v(\phi(x))-Jv(x)=0,\quad\forall u,v\in\mathscr{F}$
@@
The equivalence between the statements indicates that an explicit structure such as 21 ii) can be retrieved from “sending $\mathscr{F}$-Dirac masses onto $\mathcal{G}$-Dirac masses” as required by 21 i) and that this is rooted in a symmetrical property between $J$ and $J^{-1}$ with some infimum being attained as in 21 iii). However all three properties depend on $J$ and we would like a general characterization of (max,+)-isomorphisms, for instance showing that they are all of the form (23). This clearly will require some assumptions on $\mathscr{X},\mathscr{Y},\mathcal{G},\mathscr{F}$ as seen from 22 exhibiting an isomorphism not satisfying (23).
@@
We shall denote by $\operatorname*{Rg}_{\mathbb{R}}(A)$ the “finite” part of $\operatorname*{Rg}(A)$, i.e., $\operatorname*{Rg}_{\mathbb{R}}(A)=\operatorname*{Rg}(A)\cap\mathbb{R}^{3}=\{\sup_{1\leq j\leq 3}A(\cdot,j)+\lambda_{j}\mid\lambda_{j}\in\mathbb{R}\}$, and similarly for $B,C,D$. Observe that $\operatorname*{Rg}(A)$ is the union of $\operatorname*{Rg}_{\mathbb{R}}(A)$ and of the two vectors that are either identically $\infty$- ∞ or +$\infty$+ ∞, and similarly for $B,C,D$, so to check the isomorphim of spaces of the form $\operatorname*{Rg}(A)$, we shall limit our attention to their “finite parts”.
[…]
@@
However we claim that there is no (max,+)-isomorphism from $\operatorname*{Rg}(A)$ to $\operatorname*{Rg}(B)$ which sends the rays generated by the columns of $A$ to the rays generated by those of $B$, i.e. satisfying 21 i) To show this, we shall use the fact that a (max,+)-isomorphism must be an isometry in Hilbert’s seminorm. Recall that the Hilbert’s seminorm on $\mathbb{R}^{n}$ is defined by
@@
This example is illustrated in Figure 1. One can vizualize here that $\operatorname*{Rg}_{\mathbb{R}}(A)$ and $\operatorname*{Rg}_{\mathbb{R}}(B)$ have the same diameters in Hilbert’s seminorm (the diameter is achieved by the distance between the two first columns, in each case). It is not difficult to check that the diameter is the only invariant for (max,+) spaces with two finite generators belonging to some $\mathbb{R}^{n}$, meaning that two such spaces are isomorphic iff they have the same diameter, which explains how this example was constructed. Note also that $A_{\cdot 1},A_{\cdot 2}$ are sup-irreducible and relatively-inf-irreducible, but that $A_{\cdot 3}=\inf^{\operatorname*{Rg}(A)}((1,0,1),(0,1,1))=\sup(A_{\cdot 1},A_{\cdot 2})$ is neither, even though $\mathbb{R}e+\{A_{\cdot 1},A_{\cdot 2},A_{\cdot 3}\}$ are the only inf-irreducible elements in $\operatorname*{Rg}(A)$. This underlines that the notions of irreducibility are distinct.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. ¯$\overline{\mathbb{R}}^{\mathscr{Y}}$) for some sets $\mathscr{X}$ and $\mathscr{Y}$, with both $\mathscr{F}$ and $\mathcal{G}$ satisfying 1. Let $J$ be a (max,+)-isomorphism from $\mathscr{F}$ onto $\mathcal{G}$. Assume that $\{\delta^{\top}_{y}\}_{y\in\mathscr{Y}}\subset\mathscr{F}$ and $\{\delta^{\top}_{x}\}_{x\in\mathscr{X}}\subset\mathcal{G}$. Then the statements of 21 hold, in particular there exists $g:\mathscr{X}\to\mathbb{R}$ and a bijective $\phi:\mathscr{X}\to\mathscr{Y}$ such that $Jf=g+f\circ\phi$.

### v3

@@ 2.2 The $e_{x}$, the Dirac-like inf-irreducible functions of $\mathcal{G}$
We now turn to our candidates for inf-irreducible elements of a complete subspace $\mathcal{G}$ of $\overline{\mathbb{R}}^{\mathscr{X}}$. Define the Archimedean class of a function $f\in\mathcal{G}$ as
$f]:=\{g\in\mathcal{G}\,|\,\exists\alpha\,\in\mathbb{R},\,f-\alpha\leq g\leq f+\alpha\}.$ (17)
Let us put an order on Archimedean classes, saying that $f]\leq[g]$ if there exists $\alpha\in\mathbb{R}$ such that $f\leq g+\alpha$. We say that $[f]$ is maximal if $f]\leq[g]\implies[f]\geq[g]$.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a complete subspace of $\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. $\overline{\mathbb{R}}^{\mathscr{Y}}$) for some sets $\mathscr{X}$ and $\mathscr{Y}$. Let $J$ be a (max,+)-isomorphism from $\mathscr{F}$ onto $\mathcal{G}$. If $f\in\mathscr{F}$ is such that $[f]$ is maximal, then $[Jf]$ is also maximal.
@@
Take $g_{0}\in\mathcal{G}$ and $\alpha\in\mathbb{R}$ such that $Jf\leq g_{0}+\alpha$. Fix $f_{0}\in\mathscr{F}$ such that $g_{0}=Jf_{0}$, then $f\leq f_{0}+\alpha$. Since $[f]$ is maximal, there exists $\alpha^{\prime}\in\mathbb{R}$, $f\geq f_{0}+\alpha^{\prime}$, composing with $J$, we obtain that $Jf\geq g_{0}+\alpha^{\prime}$, so $Jf]=[g_{0}]$ and $[Jf]$ is maximal. ∎
It is thus tempting to focus on maximal Archimedean classes, and we will do so in ˜23. However some sets have no maximal classes as seen in Remark 18.
We will often need in the sequel to consider at least some finite values, hence we restrict ourselves to proper sets $\mathcal{G}$ of functions of $\overline{\mathbb{R}}^{\mathscr{X}}$. We are also not interested in situations where points cannot be distinguished by $\mathcal{G}$, since we could then quotient $\mathscr{X}$ by a $\mathcal{G}$-induced equivalence relation. Hence we introduce
@@
Let $\mathcal{G}$ be a complete subspace of $\overline{\mathbb{R}}^{\mathscr{X}}$ for some set $\mathscr{X}$, with $\mathcal{G}$ satisfying ˜1. Define, for any $x\in\mathscr{X}$, the function $e_{x}:\mathscr{X}\rightarrow\overline{\mathbb{R}}$ by
@@
with -$\infty$ absorbing. Moreover, for any $f\in\mathcal{G}$, we have the representation (with $\infty$ absorbing)
@@
and $e_{x}$ satisfies a triangle inequality: for all $x,x^{\prime},x^{\prime\prime}\in\mathscr{X}$, $e_{x}(x^{\prime\prime})\leq e_{x^{\prime}}(x^{\prime\prime})+e_{x}(x^{\prime})$. If $f\in\mathcal{G}$ is such that $[f]$ is maximal, then for all $x_{0}\in\operatorname*{Dom}(f)$ we have $e_{x_{0}}]=[f]$, i.e. we can fix $\lambda_{0}\in\mathbb{R}$, such that $e_{x_{0}}+\lambda_{0}\leq f$. If $\mathcal{G}$ contains the constant functions, then $e_{x}$ has nonnegative values.
@@
Let $x\in\mathscr{X}$. Since $\mathcal{G}$ is proper, we can fix $u_{x}\in\mathcal{G}$ such that $u_{x}(x)\in\mathbb{R}$. So $e_{x}\geq u_{x}-u_{x}(x)$, whence $e_{x}(x)\geq 0$, while by definition $e_{x}(x)\leq 0$. We have that $e_{x}\in\mathcal{G}$ as $\mathcal{G}$ is sup-stable. If $e_{x}=\inf(f,g)$ for some $f,g\in\mathcal{G}$ then w.l.o.g. we can assume that $f(x)=e_{x}(x)=0$ so, by maximality of $e_{x}$, $f\leq e_{x}=\inf(f,g)$ and $f=e_{x}$. Take $f\in\mathcal{G}$ with $[f]$ maximal. Since $f\leq e_{x}+f(x)$ for all $x\in\operatorname*{Dom}(f)$ i.e. $e_{x}]\geq[f]$, we have that $e_{x}]\leq[f]$ implies that $e_{x}]=[f]$. Fix $x\neq x^{\prime}$ and $\lambda\in\overline{\mathbb{R}}$ and $g_{1},g_{2}\in\mathcal{G}$ as in ˜1. Assume that $e_{x}=e_{x^{\prime}}+\lambda$. We can already exclude $\lambda\in\{\pm\infty\}$ since $e_{x}$ has a finite value in $x$. Moreover by definition of $e_{x}$ we have
@@
Evaluating in $x^{\prime}$ and $x$ the above expressions gives $\lambda\geq g_{1}(x^{\prime})-g_{1}(x)\geq\lambda$. The same holds for $g_{2}$, which contradicts ˜1. So $e_{x}\neq e_{x^{\prime}}+\lambda$. Eq.(19) is just a rewriting of (18).
Concerning (20), on the one hand, we can take the infimum over $x^{\prime}\in\mathscr{X}$ of $f\leq e_{x^{\prime}}+f(x^{\prime})$, the latter giving the triangle inequality when taking $f=e_{x}$. On the other hand, for any $y\in\mathscr{X}$, we have that $\inf_{x\in\operatorname*{Dom}(f)}e_{x}(y)+f(x)\leq e_{y}(y)+f(y)=f(y)$. ∎
@@
Consider the set $\mathcal{G}$ of increasing functions from $\mathbb{R}$ to $\overline{\mathbb{R}}$. It is sup and inf-stable. A quick computation gives $e_{x}=\delta^{\top}_{\{y\in\mathbb{R}|y\leq x\}}$ where $\delta^{\top}_{A}$ denotes the indicator function of a subset $A\subset\mathbb{R}$. This $e_{x}$ looks like a Dirac function, but $e_{x}]$ is strictly smaller than $e_{x^{\prime}}]$ for any $x^{\prime}<x$, so no Archimedean class is maximal. Moreover the constant functions are inf-irreducible, but do not belong to $\{e_{x}+\lambda\}_{\lambda\in\mathbb{R},x\in\mathbb{R}}$.
@@
Consider the inf-closure $\overline{\mathcal{G}}^{\inf}$ of a complete subspace $\mathcal{G}\subset\overline{\mathbb{R}}^{\mathscr{X}}$ satisfying ˜1. Then any $f\in\mathcal{G}$ such that $f$ is inf-irreducible in $\overline{\mathcal{G}}^{\inf}$ is inf-irreducible in $\mathcal{G}$ and
@@
Set $\tilde{e}_{x}^{<}(\cdot)=\sup\{u\in\overline{\mathcal{G}}^{\inf}\,|\,u(x)<0\}$ and $\tilde{e}_{x}(\cdot)=\sup\{u\in\overline{\mathcal{G}}^{\inf}\,|\,u(x)\leq 0\}$. By definition, $\tilde{e}_{x}^{<}\leq\tilde{e}_{x}$, and, for all $\epsilon>0$, $\tilde{e}_{x}-\epsilon\leq\tilde{e}_{x}^{<}$. Taking the limit $\epsilon\to 0^{+}$, $\tilde{e}_{x}^{<}=\tilde{e}_{x}$. Since $\mathcal{G}\subset\overline{\mathcal{G}}^{\inf}$, $e_{x}\leq\tilde{e}_{x}$. Let $u=\inf_{\alpha\in\mathcal{A}}h_{\alpha}\in\overline{\mathcal{G}}^{\inf}$ such that $u(x)<0$ and fix $\alpha\in\mathcal{A}$ such that $h_{\alpha}(x)\leq 0$. As $h_{\alpha}\in\mathcal{G}$, we have $u\leq h_{\alpha}\leq e_{x}$. Taking the supremum over such $u$ implies that $\tilde{e}_{x}^{<}\leq e_{x}$, which concludes the proof. ∎
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a complete subspace of $\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. $\overline{\mathbb{R}}^{\mathscr{Y}}$) for some sets $\mathscr{X}$ and $\mathscr{Y}$, with both $\mathscr{F}$ and $\mathcal{G}$ satisfying ˜1. Let $J$ be a (max,+)-isomorphism from $\mathscr{F}$ onto $\mathcal{G}$ and define $e_{x}\in\mathcal{G}$ (resp. $e^{\prime}_{y}\in\mathscr{F}$) as in (18). Then the following statements are equivalent:
@@
    for all $y\in\mathscr{Y}$, $\lambda^{\prime}\in\overline{\mathbb{R}}$, $J(e^{\prime}_{y}+\lambda^{\prime})\in\{e_{x}+\lambda\}_{x\in\mathscr{X},\,\lambda\in\overline{\mathbb{R}}}$;
@@
Fix $y\in\mathscr{Y}$ and recall that $e^{\prime}_{y}(\cdot):=\sup\{v\in\mathscr{F}\,|\,v(y)\leq 0\}$. By assumption i), $Je^{\prime}_{y}=e_{\psi(y)}+\bar{g}(y)$ with $\bar{g}(y)=Je^{\prime}_{y}(\psi(y))\in\mathbb{R}$. Conversely there exists a $\varphi:\mathscr{X}\to\mathscr{Y}$ and $\lambda_{x}\in\mathbb{R}$ such that $e_{x}=J(e^{\prime}_{\varphi(x)}+\lambda_{x})=e_{\psi(\varphi(x))}+\lambda_{x}+\bar{g}(\varphi(x))$. However from ˜17, we know that ˜1 then ensures that $x=\psi(\varphi(x))$, so $\psi$ is surjective. Similarly, consider $Je^{\prime}_{y}=Je^{\prime}_{y^{\prime}}$, to get that $\psi$ is injective.
@@
so i) holds. Assume that $Jf=g+f\circ\phi=\tilde{g}+f\circ\tilde{\phi}$ for some alternative $\tilde{g}$ and $\tilde{\phi}$ satisfying (23). Then the above computation gives that $g(\phi^{-1}(y))+e_{\phi^{-1}(y)}=\tilde{g}(\tilde{\phi}^{-1}(y))+e_{\tilde{\phi}^{-1}(y)}$ for all $y\in\mathscr{Y}$. ˜17 then gives that necessarily $\phi=\tilde{\phi}$, and, evaluating in $\phi^{-1}(y)$, $g=\tilde{g}$. So the representation (23) is unique.
@@
$0=\inf_{y\in\mathscr{Y}}Je^{\prime}_{y}(x)+J^{-1}e_{x}(y)=\inf_{y\in\mathscr{Y}}\sup_{u,v\in\mathscr{F}}Ju(x)-u(y)+v(y)-Jv(x),$
we see that, for $y=\phi(x)$, we have $0=\sup_{u,v\in\mathscr{F}}Ju(x)-u(y)+v(y)-Jv(x)$. Hence we obtain $Ju(x)-u(y)+v(y)-Jv(x)\leq 0$ for all $u,v\in\mathscr{F}$. But we also have $Jv(x)-v(y)+u(y)-Ju(x)\leq 0$. The sum of these nonpositive terms is null, so they are both null, i.e.
$Ju(x)-u(\phi(x))+v(\phi(x))-Jv(x)=0,\quad\forall u,v\in\mathscr{F}.$
@@
The equivalence between the statements indicates that an explicit structure such as ˜21 ii) can be retrieved from “sending $\mathscr{F}$-Dirac masses onto $\mathcal{G}$-Dirac masses” as required by ˜21 i) and that this is rooted in a symmetrical property between $J$ and $J^{-1}$ with some infimum being attained as in ˜21 iii). However all three properties depend on $J$ and we would like a general characterization of (max,+)-isomorphisms, for instance showing that they are all of the form (23). This clearly will require some assumptions on $\mathscr{X},\mathscr{Y},\mathcal{G},\mathscr{F}$ as seen from Remark 22 exhibiting an isomorphism not satisfying (23).
@@
I.e., we identify the matrix $A=(a_{ij})$ to the kernel $a:\{1,2,3\}^{2}\to\mathbb{R}\cup\{-\infty\}$, $(i,j)\mapsto a_{ij}$, and make similar identifications for the matrices $B,C,D$. We shall denote by $\operatorname*{Rg}_{\mathbb{R}}(A)$ the “finite” part of $\operatorname*{Rg}(A)$, i.e., $\operatorname*{Rg}_{\mathbb{R}}(A)=\operatorname*{Rg}(A)\cap\mathbb{R}^{3}=\{\sup_{1\leq j\leq 3}A(\cdot,j)+\lambda_{j}\mid\lambda_{j}\in\mathbb{R}\}$. Observe that $\operatorname*{Rg}(A)$ is the union of $\operatorname*{Rg}_{\mathbb{R}}(A)$ and of the two vectors that are either identically -$\infty$ or $\infty$. Hence to check the isomorphism of spaces of the form $\operatorname*{Rg}(A)$, we shall limit our attention to their “finite parts”.
$x_{1}$$x_{2}$$x_{3}$$A_{\cdot 1}$$A_{\cdot 3}$$A_{\cdot 2}$
$x_{1}$$x_{2}$$x_{3}$$B_{\cdot 1}$$B_{\cdot 3}$$B_{\cdot 2}$
@@
However we claim that there is no (max,+)-isomorphism from $\operatorname*{Rg}(A)$ to $\operatorname*{Rg}(B)$ which sends the rays generated by the columns of $A$ to the rays generated by those of $B$, i.e. satisfying ˜21 i) To show this, we shall use the fact that a (max,+)-isomorphism must be an isometry in Hilbert’s seminorm. Recall that the Hilbert’s seminorm on $\mathbb{R}^{n}$ is defined by
@@
This example is illustrated in Figure˜1. Observe that $\operatorname*{Rg}_{\mathbb{R}}(A)$ is a subset of $\mathbb{R}^{3}$ which is invariant by the translation by the constant vector $(1,1,1)$. So to vizualize it, it suffices to draw the cross-section of $\operatorname*{Rg}_{\mathbb{R}}(A)$ by the hyperplane orthogonal to the vector $(1,1,1)$. This cross-section is shown on the figure. One can see that $\operatorname*{Rg}_{\mathbb{R}}(A)$ and $\operatorname*{Rg}_{\mathbb{R}}(B)$ have the same diameters in Hilbert’s seminorm (the diameter is achieved by the distance between the two first columns, in each case). It is not difficult to check that the diameter is the only invariant for (max,+) spaces with two finite generators belonging to some $\mathbb{R}^{n}$, meaning that two such spaces are isomorphic iff they have the same diameter, which explains how this example was constructed. Note also that $A_{\cdot 1},A_{\cdot 2}$ are sup-irreducible and relatively-inf-irreducible, but that $A_{\cdot 3}=\inf^{\operatorname*{Rg}(A)}((1,0,1),(0,1,1))=\sup(A_{\cdot 1},A_{\cdot 2})$ is neither, even though $\mathbb{R}e+\{A_{\cdot 1},A_{\cdot 2},A_{\cdot 3}\}$ are the only inf-irreducible elements in $\operatorname*{Rg}(A)$. This underlines that the notions of irreducibility are distinct.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be a complete subspace of $\overline{\mathbb{R}}^{\mathscr{X}}$ (resp. $\overline{\mathbb{R}}^{\mathscr{Y}}$) for some sets $\mathscr{X}$ and $\mathscr{Y}$, with both $\mathscr{F}$ and $\mathcal{G}$ satisfying ˜1. Let $J$ be a (max,+)-isomorphism from $\mathscr{F}$ onto $\mathcal{G}$. Assume that $\{\delta^{\top}_{y}\}_{y\in\mathscr{Y}}\subset\mathscr{F}$ and $\{\delta^{\top}_{x}\}_{x\in\mathscr{X}}\subset\mathcal{G}$. Then the statements of ˜21 hold, in particular there exists $g:\mathscr{X}\to\mathbb{R}$ and a bijective $\phi:\mathscr{X}\to\mathscr{Y}$ such that $Jf=g+f\circ\phi$.